OpenStudy (anonymous):

find the exact value of cos75 degrees

OpenStudy (anonymous):

Use a half angle identity (150/2 = 75) or sum and difference identity (30 + 45) = 75 half angle (Θ = 150°): \[\cos \frac{ \theta }{ 2 }=\sqrt{\frac{ 1+\cos \theta }{ 2 }}\] sum and difference (u = 30°, v = 45°) \[\cos (u + v) = \cos u \cos v - \sin u \sin v\]

OpenStudy (anonymous):

That is just as confusing as the notes

OpenStudy (anonymous):

this is the example I have in the notes

OpenStudy (anonymous):

ok so it looks like they used a sum and difference identity: \(cos(u+v)= \cos u \cos v− \sin u \sin v\) The first thing you have to do is find 2 numbers on your unit circle that add up to 75. That's where the 30 and 45 come from. Then let u be 30 and let v be 45 and we're going to plug them in \(cos 75°=cos(30°+45°) With me so far?

OpenStudy (anonymous):

so far, yes lol

OpenStudy (anonymous):

I have one as a tab, not printed out

OpenStudy (anonymous):

that's good enough.

OpenStudy (anonymous):

So this is our identity cos (u + v) = cos u cos v - sin u sin v Plug in the 30 and 45 cos (30° + 45°) = cos 30° cos 45° - sin 30° sin 45° Now use the unit circle to find the 4 values in the identity: cos 30° cos 45° sin 30° sin 45°

OpenStudy (anonymous):

why do you subtract it by sin ?

OpenStudy (anonymous):

your formula is different than the one in my notes :/

OpenStudy (anonymous):

that's just the identity. Here's a list of the identities. Look under sum and difference http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

so how do you know which one to use ?

OpenStudy (anonymous):

which sign to use, or which identity?

OpenStudy (anonymous):

identity

OpenStudy (anonymous):

It depends on what you're trying to find. For this one you can use multiple identities. I started off wanting to use a half angle because then all you'd have to do is find cos 150° and plug it into the formula. You could also have use the difference identity for cosine because 120 - 45 = 75. You're just looking for any combination that works. They'll all give you the same answer

OpenStudy (anonymous):

so pretty much you can us any, just depending if you're adding or subtracting ?

OpenStudy (anonymous):

yeah as long as it's the correct trig function and the numbers add/subtract right

OpenStudy (anonymous):

okay , I understand that part so far

OpenStudy (anonymous):

cos (30° + 45°) = cos 30° cos 45° - sin 30° sin 45° ok so now we plug in the numbers from the unit circle \[\cos 75°=\frac{ \sqrt3 }{ 2 }\times \frac{ \sqrt2 }{ 2 }-\frac{ 1 }{ 2 }\times \frac{ \sqrt2 }{ 2 }\]

OpenStudy (anonymous):

\[\cos 75°=\frac{ \sqrt6 }{ 4 }-\frac{ \sqrt2 }{ 4 }\] \[\cos 75°=\frac{ \sqrt6 -\sqrt2 }{ 4 }\]

OpenStudy (anonymous):

where did the 1/2 go

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

you're multiplying xD

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

I don't remember how to add and subtract radicals. I always mix it up

OpenStudy (anonymous):

the numbers under the radicals have to be the same to add or subtract, so this can't be simplified any further

OpenStudy (anonymous):

so it'd have to be \[\sqrt{4} - \sqrt{4} \] to subtract ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but wouldn't that just always equal 0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but if you had something like this, just subtract the "coefficients" \[5\sqrt6-3\sqrt6=2\sqrt6\]

OpenStudy (anonymous):

oooh okay, it's coming back lol. It's been a longer summer

OpenStudy (anonymous):

haha everyone's saying that

OpenStudy (anonymous):

I've been taking summer school for algebra 2 and trigonometry was thrown into algebra 2, which makes no sense to me

OpenStudy (anonymous):

yeah that doesn't make sense. This is more advanced trig so it really shouldn't be mixed in there

OpenStudy (anonymous):

I passed both semester of algebra 2 with a D, that being an accelerated class . I'm taking regular algebra 2 for summer school . Idk how the state thinks this is possible for an algebra 2 class lol

OpenStudy (anonymous):

That sounds like GA. I've tutored here since 2009 and the curriculum has changed 4 times. They keep moving stuff from one class to the next. It's ridiculous

OpenStudy (anonymous):

Common core is just awful . This year is the first year for common core to be fully enforced by math . Trig is introduced in accelerated algebra 2 when I was taking the class , but it wasn't so in depth

OpenStudy (anonymous):

@peachpi could you help me with tan105 degrees ?

OpenStudy (anonymous):

105 = 150 - 45

OpenStudy (anonymous):

\[\tan(u-v)=\frac{ \tan u - \tan v }{ 1+\tan u \tan v }\] \[\tan(150°-45°)=\frac{ \tan 150° - \tan 45° }{ 1+\tan 150° \tan 45° }\]

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

to find tan , you have to divide sin by cos right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then you multiply the denominator out to only have a numerator ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okay , ima try it and see what I come up with

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you divide by the reciprocal right ?

OpenStudy (anonymous):

multiply *

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so for tan150 , I got \[-\frac{ 2 }{ 2\sqrt{3} }\] is that right ?

OpenStudy (anonymous):

uh , no . where did that come from ?

OpenStudy (anonymous):

oh, wait never mind. I misread what you had. That is correct for tan 150. I read tan 105

OpenStudy (anonymous):

oh no lol. I'm doing it step by step

OpenStudy (anonymous):

so then tan45 is 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and to substitute , I got \[\frac{ - \frac{ 2 }{ 2\sqrt{3} } - 1 }{ 1+(-\frac{ 2 }{ 2\sqrt{3} }) \times 1}\]

OpenStudy (anonymous):

Yes that's right

OpenStudy (anonymous):

how do you factor that lol

OpenStudy (anonymous):

start with simplifying: -2/(2√3) = -1/√3 Then multiply by √3/√3 to clear denominators